OpenStudy (anonymous):

HELP!!! MEDAL WILL BE GIVEN!!! Find an equation of the plane containing point B(-1,3,7) with normal n parallel to line 2 : x=2+2t, y=3t, z=-10+5t.

OpenStudy (amoodarya):

 so n=||(2,3,5) now (x-xb,y-yb,z-zb) .(2,3,5)=0

OpenStudy (amoodarya):

can you go on ?

OpenStudy (amoodarya):

(x-(-1) ,y-3,z-7) .(2,3,5) =0 2x+2 +3y -9 +5z-35=0

OpenStudy (anonymous):

Great thanks been stuck on that one for a while.